Circular No. 6576

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970228 E. Costa, M. Feroci, L. Piro, and M. N. Cinti, Istituto di Astrofisica

Spaziale (IAS), CNR, Frascati; F. Frontera and G. Zavattini, Universitá di Ferrara; L. Nicastro, E. Palazzi, D. Dal Fiume, and M. Orlandini, Istituto Tecnologie e Studio Radiazioni Extraterrestri (ITESRE), CNR, Bologna; J. in ’t Zand, J. Heise, and R. Jager, Space Research Organization of the Netherlands (SRON), Utrecht; A. Parmar and A. Owens, Space Science Department, ESTEC; S. Molendi, Istituto di Fisica Cosmica e Tecnologie Relative, CNR, Milano; G. Cusumano, M. C. Maccarone, and S. Giarrusso, Istituto di Fisica Cosmica e Applicazioni Informatica, CNR, Palermo; L. A. Antonelli, F. Fiore, P. Giommi, J. M. Muller, L. Salotti, G. Gennaro, M. Stornelli, G. Crisigiovanni, R. Ricci, and A. Coletta, BeppoSAX, Rome; and R. C. Butler, Agenzia Spaziale Italiana, Rome, report: “The BeppoSAX GRBM/WFC error box of GRB 970228 (IAUC 6572) was observed with the BeppoSAX Narrow Field Instruments on Feb. 28.458 UT for 15 000 s (8 hr after the γ-ray burst). A previously unknown intense x-ray source, SAX J0501.7+1146, has been detected by the MECS and LECS at the same position (α = 5^{h}01^{m}44^{s}, δ = +11^{o}46′.7, equinox 2000.0; estimated error radius 50′′). This position lies at the edge of the reported BeppoSAX WFC error box (IAUC 6572). The source flux is (2.8 ± 0.4) × 10^{−12} erg cm^{−2} s^{−1} in the MECS (2–10 keV) and (4.0 ± 0.6) × 10^{−12} erg cm^{−2} s^{−1} in the LECS (0.5–10 keV). The field was observed again on Mar. 3.734, and a source was detected at a position consistent with the previous one, but at a flux level lower by a factor of 20. Given this very peculiar time variability, we conclude that the new BeppoSAX x-ray source is related to GRB 970228.”

D. A. Frail, National Radio Astronomy Observatory; S. R. Kulkarni, California Institute of Technology; E. Costa, M. Feroci, and L. Piro, IAS, Frascati; L. Nicastro, D. Dal Fiume, and E. Palazzi, ITESRE, Bologna; F. Frontera, Universitá di Ferrara; and R. Jager and J. Heise, SRON, report: “Observations at 20 cm with the Very Large Array on Mar. 1.03 and 2.16 UT, and at 6 cm on Mar. 4.04, centered on the γ-ray burst position given on IAUC 6572, have detected a single, unresolved radio source in the error circle of the x-ray transient seen by Costa et al. (above). The radio source is at α = 5^{h}01^{m}41^{s}.68 ± 0^{s}.01, δ = +11^{o}47′05′′.1 ± 0′′.2 (equinox J2000.0). Flux densities measured on these three days are 0.80, 0.76, and 0.62 mJy, respectively, with errors of order 0.15 mJy.”

1997 March 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green